# RefVNLI: Towards Scalable Evaluation of Subject-driven Text-to-image Generation

### v1

@@ 3.3 Baselines
We compare RefVNLI with two groups of automatic metrics for textual alignment. The first group leverages large vision-language models (VLMs), computing cosine similarity between text and image encodings. This includes BLIPScore [33], CLIPScore [15], and SigLIP [73]. The second group, which includes TIFA [19] and VQAScore [35], evaluates textual alignment via visual question answering (VQA). We also include a baseline where PaliGemma is finetuned on our dataset exclusively for textual alignment, given only the prompt and target image, referred to as PaliGemmatext.
Baselines for subject consistency.
For subject consistency, we compare RefVNLI against baselines that rely on large VLMs by measuring cosine similarity between the reference and generated image encodings. These baselines include DINO-I [7], and Crop-IR [66]. For the Human category, we also use ArcFace [9], a face-embedding model designed for identity recognition. Furthermore, similar to textual alignment, we evaluate against a PaliGemma model finetuned on our dataset specifically for subject consistency evaluation, using only the reference and target images (formatted as in Sec. 3.1). We refer to this model as PaliGemmaref.
Textual Alignment Subject Consistency Unified Evaluation
Animal Human Object Animal Human Object Animal Human Object
CLIP 72.8 77.4 74.6 72.4 87.7 76.4 72.6 82.2 75.5
DINO - - - 80.1 78.0 77.4 - - -
Crop-IR - - - 56.1 56.7 53.4 - - -
ArcFace - - - - 61.1 - - - -
CLIPScore 71.5 76.2 72.9 - - - - - -
BLIPScore 75.4 79.6 78.8 - - - - - -
SigLIP 72.5 80.1 77.1 - - - - - -
TIFA 70.6 75.7 69.5 - - - - - -
VQAScore 79.3 78.1 82.6 - - - - - -
VIEScore 76.6 78.1 78.5 65.4 80.3 76.0 70.6 79.2 77.2
DreamBench++ 79.5 82.7 82.5 74.5 84.2 79.4 76.9 83.4 80.9
@@
RefVNLI 80.1 82.5 82.0 79.4 86.0 85.7 79.7 84.2 83.8
Table 1: ROC AUC scores on DreamBench++ for textual alignment, subject consistency, and their harmonic mean (as a unified evaluation) across Animal, Human, and Object categories. The last two rows feature models finetuned on our dataset, with PaliGemmatext/ref comprising two separate models (PaliGemmatext and PaliGemmaref) trained exclusively for each criterion. Bold indicates the highest score per column.
Textual Alignment Subject Consistency Unified Evaluation
CLIP 81.7 74.7 81.1 63.9 73.3 52.6 71.7 74.0 63.8
DINO - - - 81.6 77.4 50.1 - - -
Crop-IR - - - 62.0 54.5 51.3 - - -
CLIPScore 81.7 74.9 79.2 - - - - - -
BLIPScore 82.9 79.7 84.2 - - - - - -
SigLIP 80.6 80.5 82.3 - - - - - -
TIFA 79.7 76.1 79.2 - - - - - -
VQAScore 77.4 83.8 87.8 - - - - - -
VIEScore 62.1 54.0 71.6 56.5 49.4 50.1 59.2 51.6 59.0
DreamBench++ 86.4 85.5 88.2 71.0 84.0 54.3 77.9 84.7 67.2
@@
RefVNLI 84.4 89.5 86.3 80.4 83.8 62.8 82.4 86.6 72.7
Table 2: ROC AUC scores on ImagenHub for textual alignment, subject consistency, and their harmonic mean (as a unified evaluation) across Animal and Object categories, as well as for the Multi-subject setting. The last two rows feature models finetuned on our dataset, with PaliGemmatext/ref comprising two separate models trained exclusively for each criterion. Bold indicates the highest score per column.
Baselines for both criteria.
We also include 3 metrics that assess both textual alignment and subject preservation. CLIP [47] computes scores separately for each criterion by calculating cosine similarity between the encodings of imagetgt and those of prompt and imageref. VIEScore [27] uses an elaborate GPT-4o [42] few-shot strategy, simultaneously generating two 0–10 ratings, one for each criterion. Finally, DreamBench++ [46] evaluates each criterion separately using distinct GPT-4o prompts with hand-crafted instructions and examples. This method follows a two-step prompting process, where GPT-4o first summarizes the evaluation task to increase task comprehension before assigning a 0–4 score.
Textual Alignment Subject Consistency Unified Evaluation
CLIP 83.3 80.2 81.7
DINO - 85.2 -
Crop-IR - 85.5 -
CLIPScore 83.4 - -
BLIPScore 82.5 - -
SigLIP 75.5 - -
TIFA 90.6 - -
VQAScore 88.9 - -
VIEScore 82.6 87.4 84.9
DreamBench++ 87.0 89.8 88.4
@@
RefVNLI 97.0 82.3 89.0
Table 3: ROC AUC scores on KITTEN (landmarks) for textual alignment, subject consistency, and their harmonic mean (as a unified evaluation). The last two rows feature models finetuned on our dataset, with PaliGemmatext/ref consisting of separate models for each criterion. Bold indicates the highest score per column.
4 Results
Our main results are summarized in Tab. 1, Tab. 2, and Tab. 3, with qualitative examples in Fig. 2.
For DreamBench++ (Tab. 1), RefVNLI ranks among the top models for textual alignment, either outperforming all or slightly trailing larger models. For subject consistency, it is among the top-2 models across all categories, and performs best in the Object category, surpassing the next-best GPT-4o-based DreamBench++ by 6.3 points.
For ImagenHub (Tab. 2), RefVNLI ranks in the top-2 for textual alignment over Animals and achieves the highest score in Objects, surpassing the best non-finetuned model by 4 points. It also excels in the Multi-subject setting, ranking top-3. In subject consistency, RefVNLI continues to rank among the top-2 models for both Animal and Object categories, while performing best in the Multi-subject setting.
Finally, for KITTEN (Tab. 3), RefVNLI achieves the highest textual alignment score but struggles with subject consistency. We hypothesize this may be due to its identity-sensitive training, which penalizes even minor mismatches in identity-defining traits (§2.1). Given that landmarks often have intricate details, this may have led to hypersensitivity to minor discrepancies, as can be observed in Fig. 11.
Textual Alignment Subject Consistency Unified Evaluation
RefVNLI (ours) 81.5 87.7 / 86.3 97 82.7 83 / 62.8 82.3 82.1 85.3 / 72.7 89.0
      reverse classification order 80 85.2 / 85.5 95.3 80.9 84.3 / 68.7 87 80.4 84.7 / 76.2 91.0
      multiclass 79.5 83.7 / 84.7 94.7 79.6 76 / 61.1 86.3 79.5 79.7 / 71 90.3
      no markup 78.4 87 / 84.3 92.3 65.5 75.9 / 60.8 88.7 71.4 81.1 / 70.6 90.5
Table 4: Ablation Study: ROC AUC scores for various ablated versions of RefVNLI across benchmarks (over all subjects). The ablations test alternative output formulations, including reversing classification order, using a four-label multiclass framework, and incorporating a designated token for separate aspect classification. We also examine a variant that excludes the subject markup from the input caption, as well as one that concatenates the reference and target images into a single image input.
Textual Alignment Visual Quality Overall Preference
CLIP 51.9 91.3 69.3
DINO - 91.3 -
Crop-IR - 87.1 -
CLIPScore 47.3 - -
BLIPScore 39.4 - -
SigLIP 74.6 - -
VQAScore 52.3 - -
VNLI - -
VIEScore 60.6 65.1 69.3
DreamBench++ 56.4 82.9 78.8
@@ 4 Results
RefVNLI 87.1 95.5 91.3
Table 5: Results on the ImageRAG rare concepts benchmark, where users compare image pairs and select the better one based on textual alignment, visual quality, and overall preference. We report accuracy in predicting these human choices, with the overall preference score computed as the harmonic mean of the textual and visual scores. The last two rows include models finetuned on our dataset, with PaliGemmatext/ref consisting of separate models for each criterion. Bold indicates the highest score per column.
Notably, Fine-tuning exclusively for textual alignment (PaliGemmatext) slightly reduces performance, particularly for Animals and Humans, while training solely for subject consistency (PaliGemmaref) results in a more significant drop, with up to a 14.8-point decrease for Humans (Tab. 1). This suggests that joint training provides complementary benefits, with subject consistency gaining the most.
Qualitative examples in Fig. 2 further illustrate RefVNLI’s strengths, highlighting its sensitivity to subtle textual alignment mismatches, such as a missing waterfall (top-left) or grassy field (mid-right versus mid-center). For subject preservation, it demonstrates robustness to identity-agnostic changes, like a zoomed-out parrot (top-center) or different clothing (bottom-center), while remaining sensitive to key identity traits, such as facial features (bottom-left) and color variations (left and middle balloons).
Overall, while RefVNLI is not always the top performer, it consistently ranks among the top-3 models. More importantly, it achieves the best balance between textual alignment and subject consistency, consistently leading in the Unified Evaluation score across all benchmarks.
5 Analysis
Figure 6: ImageRAG Rare Entities Examples: We compare RefVNLI with CLIP and DreamBench++ in aligning with human preferences (top rows of each example) across Textual Alignment (TA), Image Quality (IQ), and Overall Preference (OP). DreamBench++ scores (0–4) are rescaled to 0–100 for readability. The higher of the two criterion-wise scores is emphasized unless both are equal. RefVNLI consistently aligns with human judgments across all three criteria. Notably, in the bottom example, it is the only metric to correctly identify the higher-quality image based on IQ, albeit by a small margin. This case is particularly challenging as it is out-of-distribution for RefVNLI, being that the preferred image is inspired by the reference rather than being of the same identity.
5.1 Ablation Study
To assess the impact of various design decisions in RefVNLI, we run an ablation study examining alternative input and output configurations. On the output side, we test (1) reversing the classification order (subject preservation before textual alignment), (2) a 4-label multiclass framework for joint text-image alignment classification, and (3) a model that prefixes a designated token (‘TEXT’ or ‘IMAGE’) to the prompt to enable separate classification of each aspect within a unified model. For inputs, we explore the effect of removing subject markup from the prompt and of concatenating the imageref and imagetgt instead of processing them separately.
Results (Tab. 4) show that reversing the classification order degrades performance, particularly in subject preservation, as does evaluating each aspect separately. This suggests that prioritizing textual alignment helps in subject preservation assessment. The multiclass approach also underperforms compared to our dual binary classification setup, highlighting the benefits of treating each criterion independently. Further, removing subject markup weakens subject preservation, underscoring its role in linking the reference image to the prompt. Lastly, concatenating images instead of processing them separately harms performance, emphasizing the advantage of distinct image inputs.
5.2 Applicability to Rare Entities
To assess RefVNLI on unfamiliar subjects, we use the rare entities benchmark from ImageRAG [53], which evaluates generated images based on prompts and reference images of uncommon subjects (e.g., scientific animal names, lesser-known dishes). Human annotators compared image pairs, selecting the better one based on emphTextual Alignment, Visual Quality (focusing on overall depiction rather than subject accuracy), as well as Overall Preference.
As shown in Tab. 5, RefVNLI consistently outperforms all baselines in aligning with human preferences across these criteria, showcasing strong robustness to rare subjects. [^4] TIFA was excluded due to assigning identical scores to 61% of pairs, making accuracy calculations unreliable. This is further supported by Fig. 6, where only RefVNLI repeatedly matches human selections. Notably, in the bottom example, it is the only metric to correctly identify the better option in terms of Visual Quality, despite the challenge of this out-of-distribution case where the superior option is inspired by, rather than preserves the identity of, the reference.
6 Related Work
Evaluation of Visual Language Models (VLMs) spans various settings, including visual reasoning [5, 25] and visual question-answering [2, 39, 40]. For text-to-image (T2I) models, assessments normally focus on image quality [16, 52], diversity [48], and alignment with the text [15, 73, 70, 19, 35, 47]. Assessing subject preservation, which is crucial for subject-driven generation, is typically done using embedding-based metrics like CLIP [47] and DINO [7]. More comprehensive metrics, like VIEScore [27] and DreamBench++ [46], employ GPT-4o [42] to measure both textual alignment and subject consistency.
Subject-driven T2I models have gained significant traction in recent years, with some approaches fine-tuning general models into specialist versions that capture specific subjects and styles [10, 29, 43, 51, 57]. Other works enable broader applicability of subject- or style-guided image generation, via one-shot examples. These methods typically fall into 2 categories: adapter-based approaches [23, 11, 64, 71], which encode reference images for integration into the diffusion model, and adapter-free methods [14, 37, 38], which directly incorporate extracted features like attention maps.
Closely related, image editing complements subject-driven T2I generation in that the generated image’s appearance is primarily governed by the input image, with the text only impacting specific aspects (unlike our setting, where the input image only affects one aspect, with text guiding the rest). The task has evolved from pixel-to-pixel translation for predefined transformations [21, 60, 77] to more flexible, text-guided modifications [6, 44, 59], with recent diffusion-based approaches improving precision via cross-attention manipulation [13, 69]. Beyond images, personalized generation extends to other modalities, including videos and texts. Video generation can be conditioned on text [17, 34, 55], reference images [65, 76], or other videos [28]. In text generation, efforts focus on style transfer [50, 74], debiasing [49, 75], and enabling broader semantic control [54, 56, 68].
Lastly, several studies leveraged intra-frame relationships in videos to learn more human-aligned visual representations. These works aim to improve robustness to identity-agnostic variations (e.g., rotation, lighting), by analyzing consecutive frames sourced from public video datasets [24, 45, 61, 62, 67] or captured by cameras on moving agents [1, 22].
7 Conclusion
In this work, we introduced RefVNLI, a reliable, cost-effective metric for subject-driven text-to-image generation, addressing both textual alignment and subject preservation. Trained on a large-scale, auto-generated dataset, RefVNLI balances robustness to identity-agnostic variations (e.g., pose, lighting, background) with sensitivity to identity-specific traits (e.g., facial features, object shape, and unique details) when evaluating subject preservation. For textual alignment, we generate subject-specific prompts, including perturbations for hard negatives, enhancing the model’s ability to detect subtle mismatches. Evaluations on several benchmarks show that RefVNLI consistently matches or surpasses all baselines, including much larger GPT-4o-based ones, especially for less-common subjects.
Future work should enhance RefVNLI’s ability to evaluate subject preservation across artistic styles, including when the reference represents a style rather than a specific entity, as well as handle textual modifications that explicitly alter identity-defining attributes (e.g., changing an object’s color). Further improvements should also focus on processing multiple reference images, both for the same subject and for distinct subjects. Overall, by providing a cost-efficient and human-aligned evaluation framework, this work aims to drive progress in personalized image generation, fostering more precise and reliable subject-driven T2I methods.
Supplementary Material
8 Reproducibility
We fine-tuned PaliGemma [4], on a balanced version of our dataset (by undersampling the more common labels). We finetuning took 24 hours, using 2 NVIDIA A100 cards, each with 80GiB memory.

### v2

@@ 3.2 Baselines
We compare RefVNLI with two groups of automatic metrics for textual alignment. The first group leverages large vision-language models (VLMs), computing cosine similarity between text and image encodings. This includes BLIPScore Li et al. (2022), CLIPScore Hessel et al. (2021), and SigLIP Zhai et al. (2023). The second group, which includes TIFA Hu et al. (2023) and VQAScore Lin et al. (2024), evaluates textual alignment via visual question answering (VQA). We also include a baseline where PaliGemma is finetuned on our dataset exclusively for textual alignment, given only the prompt and target image, referred to as PaliGemmatext.
Baselines for subject preservation.
For subject preservation, we compare RefVNLI to baselines that use large VLMs by computing cosine similarity between reference and target image embeddings. These include DINO Caron et al. (2021), Crop-IR Winter et al. (2024), [^5] See Appendix E for more details. and for the Human category, also ArcFace Deng et al. (2019), a face-recognition model. We also assess a PaliGemma model finetuned on our dataset solely for subject preservation, using only reference and target images (formatted as in §2.2), denoted as PaliGemmaref.
Baselines for both criteria.
We also include 3 metrics that assess both criteria. CLIP Radford et al. (2021) computes scores separately for each criterion by calculating cosine similarity between the encodings of imagetgt and those of prompt and imageref. VIEScore Ku et al. (2024a) uses an elaborate GPT-4o OpenAI (2024) few-shot strategy, simultaneously generating two 0–10 ratings, one for each criterion. Lastly, DreamBench++ Peng et al. (2024) evaluates each criterion separately using distinct GPT-4o prompts with hand-crafted instructions and examples. This method follows a two-step prompting process, where GPT-4o first summarizes the evaluation task to increase task comprehension before assigning a 0–4 score.
Textual Alignment Subject Preservation Unified Evaluation
CLIP 81.8   74.7   81.1   63.8   73.3   52.6   71.6   74.0   63.8
DINO - - - 81.7   77.3   50.0 - - -
Crop-IR - - - 77.6 84.1 56.8 - - -
CLIPScore 81.5   75.0   79.1 - - - - - -
BLIPScore 82.9   79.7   84.2 - - - - - -
SigLIP 80.7   80.6   82.3 - - - - - -
TIFA 79.9   76.1   79.2 - - - - - -
VQAScore   77.3   83.8 87.8 - - - - - -
VIEScore   62.1   54.1   71.6   56.4   49.4   50.2   59.0   51.5   58.9
DreamBench++ 86.4   85.5 88.2   71.1 84.0   54.3   78.0 84.8   67.2
@@
RefVNLI 84.6 89.4 86.2 80.2 83.8 62.7 82.3 86.5 72.6
Table 2: ROC AUC scores on ImagenHub for textual alignment, subject preservation, and their harmonic mean (as a unified evaluation) across Animal and Object categories, as well as for the Multi-subject setting.
Textual Alignment Subject Preservation Unified Evaluation
CLIP   83.2 80.1   81.5
DINO - 85.4 -
Crop-IR - 90.2 -
CLIPScore   83.3 - -
BLIPScore   82.6 - -
SigLIP   75.3 - -
TIFA   90.6 - -
VQAScore   89.0 - -
VIEScore   82.5 87.5 84.9
DreamBench++   87.0   89.9 88.4
@@
RefVNLI 97.0 82.2 88.9
Table 3: ROC AUC scores on KITTEN (landmarks) for textual alignment, subject preservation, and their harmonic mean (as a unified evaluation).
4 Results
Our main results are summarized in Tables 1, 2, and 3, with qualitative examples in Fig. 2.
On DreamBench++ (Table 1), RefVNLI outperforms or statistically matches all baselines across both criteria, with a notable 6.3-point lead over the GPT-4o-based DreamBench++ metric in subject preservation for Objects. This is especially notable given the benchmark’s diverse visual styles, including cartoonish and pixelated images, which are outside RefVNLI’s training distribution of real-world video frames. Similarly, on ImagenHub (Table 2), RefVNLI matches or exceeds all baselines in both single- and multi-subject settings, with 5.9-point gains over the strongest non-finetuned model on subject preservation of the multi-subject setting (Crop-IR). Lastly, on KITTEN (Table 3), RefVNLI leads in textual alignment but underperforms in subject preservation, though it remains statistically comparable to most baselines. This may result from RefVNLI’s identity-sensitive training, which penalizes minor deviations—especially challenging for landmarks with intricate visual details (Fig. 12)—and from a domain shift, as landmarks were absent from RefVNLI’s training data (OOD).
Notably, across all benchmarks, fine-tuning only for textual alignment (PaliGemmatext) slightly reduces performance, especially for Animals and Humans, while training solely for subject preservation (PaliGemmaref) yields even larger declines—up to a 14.8 points for Humans (Table 1). This suggests that joint training provides complementary benefits, with subject preservation gaining the most.
Fig. 2 further showcases RefVNLI’s strengths, like its sensitivity to subtle textual alignment errors, such as a missing waterfall (top-left). For subject preservation, it remains robust to identity-agnostic changes, like a zoomed-out parrot or person (top-center and bottom-center) or different clothes (bottom-center), while staying sensitive to key identity traits, e.g., changed facial features (bottom-left) and colors (left and middle balloons).
Overall, RefVNLI consistently outperforms or statistically matches all baselines on both criteria, with the only exception of subject preservation in the OOD landmarks category, where it still performs competitively. Importantly, it offers the best trade-off between textual alignment and subject preservation, surpassing all non-finetuned metrics in Unified Evaluation across all benchmarks.
5 Applicability to Rare Entities
To test RefVNLI on unfamiliar subjects, we use the ImageRAG benchmark Shalev-Arkushin et al. (2025), which evaluates generated images based on prompts and reference images of uncommon subjects (e.g., scientific animal names, lesser-known dishes). Human annotators compared image pairs, selecting the better one based on Textual Alignment, Visual Quality (evaluating general depiction of the entity rather than exact reference-adherence), and Overall Preference. We report per-axis accuracy, defined as the frequency with which a metric ranks the human-preferred image higher, with approximate ties handled by rounding. Overall Preference is computed as the harmonic mean of textual and visual scores, and significance testing is assessed via bootstrap resampling, akin to §3.1.
As shown in Table 4, RefVNLI consistently outperforms all baselines in aligning with human preferences across these criteria, showcasing strong robustness to rare subjects. [^6] TIFA was excluded due to assigning identical scores to 61% of pairs, making accuracy calculations unreliable. This is further supported by Fig. 6, where only RefVNLI repeatedly matches human selections.
Textual Alignment Visual Quality Overall Preference
CLIP   51.8 91.3   69.2
DINO - 91.4 -
Crop-IR -   86.4 -
CLIPScore   47.4 - -
BLIPScore   39.6 - -
SigLIP   74.8 - -
VQAScore   52.3 - -
VIEScore   60.8   65.4   69.6
DreamBench++   56.6   83.1   78.9
@@ 5 Applicability to Rare Entities
RefVNLI 87.2 95.5 91.4
Table 4: Results on ImageRAG rare concepts, where users select the better image in each pair based on textual alignment, visual quality (general entity depiction rather than specific subject adherence), and overall preference. We report accuracy: how often models ranked the human-preferred image higher. Overall preference is the harmonic mean of textual and visual scores.
6 Related Work
Evaluation of Visual Language Models (VLMs) spans various settings, including visual reasoning Bitton-Guetta et al. (2024); Kahou et al. (2017) and visual question-answering Antol et al. (2015); Marino et al. (2019); Mensink et al. (2023). For text-to-image (T2I) models, assessments normally focus on image quality Heusel et al. (2017); Salimans et al. (2016), diversity Rassin et al. (2024), and alignment with the text Hessel et al. (2021); Radford et al. (2021); Hu et al. (2023); Yarom et al. (2023); Zhai et al. (2023); Lin et al. (2024). Assessing subject preservation, which is crucial for subject-driven generation, is typically done using embedding-based metrics like CLIP Radford et al. (2021) and DINO Caron et al. (2021). Other metrics, like VIEScore Ku et al. (2024a) and DreamBench++ Peng et al. (2024), use GPT-4o OpenAI (2024) to measure both criteria.
Subject-driven T2I models have been gaining much traction, with some methods fine-tuning general models into specialist versions that capture specific subjects and styles Gal et al. (2022); Kumari et al. (2023); Ruiz et al. (2023); Sohn et al. (2023); Park et al. (2024). Others focus on broader applicability using one-shot examples, either through adapter-based methods that integrate encoded reference images into diffusion models Gal et al. (2023); Jia et al. (2023); Wei et al. (2023); Ye et al. (2023) or via adapter-free techniques that directly use extracted features such as attention maps Liu et al. (2023); Hertz et al. (2024); Lv et al. (2024).
Closely related, image editing complements subject-driven T2I generation in that the generated image’s appearance is primarily governed by the input image, with the text only impacting specific aspects, whereas in out setting it is the other way around. The task has evolved from pixel-to-pixel translation for predefined transformations Isola et al. (2017); Zhu et al. (2017); Wang et al. (2018) to more flexible, text-guided edits Brooks et al. (2023); Tumanyan et al. (2023); Parmar et al. (2023), with recent diffusion-based methods improving precision via cross-attention manipulation Hertz et al. (2022); Yang et al. (2023). Beyond images, personalized generation extends to other modalities, including videos and texts. Video generation can be conditioned on text Li et al. (2018); Hong et al. (2022); Singer et al. (2022), reference images Wei et al. (2024); Zhou et al. (2024), or other videos Ku et al. (2024c). In text generation, efforts focus on style transfer Reif et al. (2022); Zhang et al. (2024), debiasing Zhao et al. (2018); Ravfogel et al. (2020), and broader semantic control Shapira et al. (2022); Slobodkin et al. (2023); Xie et al. (2023).
Figure 6: ImageRAG Rare Entities Examples: We compare RefVNLI with CLIP and DreamBench++ in aligning with human preferences (top rows of each example) across Textual Alignment (TA), Image Quality (IQ), and Overall Preference (OP). DreamBench++ scores (0–4) are rescaled to 0–100 for readability. The higher of the two criterion-wise scores is emphasized unless both are equal.
Finally, several studies leveraged intra-frame relationships in videos to learn more human-aligned visual representations. These works aim to improve robustness to identity-agnostic variations (e.g., rotation, lighting), by analyzing consecutive frames sourced from public video datasets Jin et al. (2018); Parthasarathy et al. (2023); Wang and Gupta (2015); Wang et al. (2017); Wu and Wang (2021) or captured by cameras on moving agents Agrawal et al. (2015); Jayaraman and Grauman (2015).
7 Conclusion
We present RefVNLI, a cost-effective and reliable metric for subject-driven T2I evaluation that jointly assesses textual alignment and subject preservation. Trained on a large-scale, auto-generated dataset, RefVNLI is designed to be robust to identity-agnostic visual variations (e.g., pose, lighting, background) while remaining sensitive to identity-specific features (e.g., facial features, object shape, and unique details) when evaluating subject preservation. For textual alignment, it leverages subject-specific prompts with perturbed hard negatives to detect and penalize fine-grained mismatches. Across benchmarks, RefVNLI outperforms or rivals all baselines, including larger GPT-4o-based metrics, particularly on less-common subjects.
Future work should focus on improving performance across artistic styles as well as identity-altering edits, and supporting multiple reference images. More broadly, RefVNLI facilitates progress in personalized T2I generation by enabling better checkpoint selection, reinforcement learning, and iterative model refinement.
8 Limitations
RefVNLI was trained on data sourced from real-life video frames and images. While it performs well on stylistically consistent inputs, including cartoonish or pixelated images, it struggles with cross-style scenarios where imageref and imagetgt differ in style, as well as when the subject undergoes explicit modifications (e.g., changes in color or shape). Additionally, the current framework is limited to single-reference cases and should be extended to support multiple references, both for the same subject and for distinct ones.
Moreover, research on subject-driven generation could benefit from a unified score capturing overall performance, rather than the two separate scores currently provided by RefVNLI. Although the harmonic mean of the two offers a reasonable proxy, future iterations should aim to output a single, integrated metric, alongside the individual, more granular scores.
Appendix A Reproducibility
@@
At inference, we used the same input structure, including the insertion of markups around the subject in the prompt and separate image inputs. The textual alignment and subject preservation scores were then computed from the probabilities of the first and second generated tokens, respectively, with each score defined as the probability of token ‘1’ divided by the sum of probabilities for tokens ‘0’ and ‘1’.
[…]
Appendix B Additional Details on the Meta-evaluation
For meta-evaluation, we perform bootstrapping (Efron, 1987) (1,000 samples per benchmark, with repetitions). For the ImageRAG rare entities benchmark, consisting of only 26 instances, we sample each time a sample 4 times this size, to ensure there are at least 100 instances. For the other benchmarks, which are significantly larger, the size of each sample is identical to the original size of the corresponding benchmark. For the calculations of confidence intervals (CIs), we use significance level $p=0.05$.